OpenStudy (anonymous):

find dy/dx in 2x^2+3xy-5y=6

OpenStudy (badhi):

There are two ways one is to take y to a one side and x terms to other side and find dy/dx as usual way The other method is just differentiate left and right side with respect to x in this method you'll have to use the chain rule (not in this case) $$2x^2+3xy−5y=6$$ $$2\frac{dx^2}{dx}+3\frac{dxy}{dx}−5\frac{dy}{dx}=0$$ $$2(2x)+3\left(y+x\frac{dy}{dx}\right)−5\frac{dy}{dx}=0$$ now take dy/dx to one side and then if you want you can substitute y with x from the given equation